Solve 4 sinx + cosecx =4
Answer by Alan3354(69443)

4 sinx + cosecx =4
4sin + 1/sin = 4
4sin^2 + 1 - 4sin = 0
4sin^2 - 4sin + 1 = 0
(2sin -1)^2 = 0
sin(x) = 1/2
x = pi/6, 5pi/6 + n*2pi, n = any integer